OpenStudy (doggy):

(i haven't been taught this lesson yet but i would like to answer this question) Which values of x will make the absolute value equation true? |x – 3| = 40 A. {–43, –37} B. {43, –37} C. {40, –40} D. {43, 37}

OpenStudy (anonymous):

To solve this, substitute values of x given in your multiple choice into the equation given.

OpenStudy (anonymous):

Example, |x-3| = 40. Let's try the x values given in A. x = , -43, and -37. Let's do x=-43 first. |(-43) - 3| = 40 |-46| = 40 46 = 40. As you can see, the answer already cannot be A because 46 does not equal 40.

OpenStudy (doggy):

ohh i see

OpenStudy (doggy):

hold on, lemme try to find the answer and i'll ask if i got it right

OpenStudy (doggy):

is it C?

OpenStudy (joannablackwelder):

That is true, @IParke , but you can only use that method effectively if you are given answer choices. Would you like to know how to solve the problem directly, @doggy?

OpenStudy (doggy):

sure, the more i learn the better i suppose

OpenStudy (anonymous):

|x – 3| = 40 is the same as saying: x - 3 = 40 x - 3 = -40 So, solve those two algebra equations. x = 43 x = -37 So B. {43, -37} is the answer.

OpenStudy (anonymous):

@JoannaBlackwelder , you are completely correct! For simplicity's sake, I am a huge fan of substitution. Fast, easy, mostly painless. :)

OpenStudy (joannablackwelder):

True, @IParke , it is a great choice, when it is an option. :)